96) Assume the desired reserve ratio is 25 percent and the Bank of Canada buys $4 million of securities from the public. As a result of this transaction the supply of money is: A) directly increased by $4 million and the money-creating potential of the chartered banking system is increased by $16 million. B) directly reduced by $4 million and the money-creating potential of the chartered banking system is decreased by $12 million. C) directly increased by $4 million and the money-creating potential of the chartered banking system is increased by $12 million. D) not directly affected, but the money-creating potential of the chartered banking system is increased by $12 million.

Answer: C

PLEASE STEP BYSTEPP SOLUTION!!
Answer #1

Total money supply = Money multiplier * Initial deposit

Money Multiplier = 1 / Reserve ration = 1 / 25% = 4

When bank of Canada buy $4 million of securities from the public then the payment is made by deposits in respective accounts.

Now deposits of bank of Canada increases by $4 million and it will increases directly the overall money supply in the market by $4 million.

Every loan is given by the bank is in form of deposits there is no cash payment of loans. Hence total money supply the $4 million deposits is ($4 * 4 ) = $16 million. And total amount of Potential money creation which can be created Bank ( Total money - reserve ) = ( $16 - $4) = $12 million

Hence $12 million dollar is the overall money creation potential of the chartered bank. Since out of $16 million of total money supply $4 million is reserved with the bank only. Actually money created is only $12 million.